# Learning Human-Perceived Fakeness in AI-Generated Videos via Multimodal LLMs

### v1

@@ 1 Introduction
Recent advances in video generation technologies, including Veo3 (Anil et al., 2024), Sora (OpenAI, 2024), Pika (Pika, 2024), Meta Movie Gen (AI, 2024), Gen-3 (Runway Research, 2024), Kling (Kling, 2024), and others (Yang et al., 2024; Team, 2024; MiniMax, 2024; Wang et al., 2023; Li et al., 2024b), have demonstrated remarkable capabilities in producing increasingly realistic videos. Alongside this progress, numerous studies about video generation have been conducted (Huang et al., 2023; Liu et al., 2023; Bansal et al., 2024; Huang et al., 2024; Liu et al., 2025), such as evaluating video prompt alignment towards a set of provided prompts as in VBench (Huang et al., 2023), or analyzing the physical commonsense in deepfake videos as explored by VideoPhy (Bansal et al., 2024), etc. However, these evaluations primarily compare AI-generated videos against a set of predetermined criteria, neglecting one of the most fundamental aspect:
@@ 4 Related Work
Text-to-Video (T2V) generation produces videos from textual prompts using Transformer and diffusion models (Vaswani et al., 2023; Ho et al., 2020). Closed-source systems (OpenAI, 2024; AI, 2024; Pika, 2024; Runway Research, 2024; Kling, 2024) showcase strong visual quality, while open-source models like Mochi and CogVideoX also achieve competitive results (Peebles & Xie, 2023; Team, 2024; Yang et al., 2024). Evaluation typically relies on vision-based scores such as IS, FID, and FVD, or multimodal benchmarks like VBench and VideoPhy (Barratt & Sharma, 2018; Heusel et al., 2018; Unterthiner et al., 2019; Huang et al., 2023; Bansal et al., 2024). However, these methods emphasize predefined attributes (e.g., object count, appearance, alignment) and overlook the human-centric question of whether viewers can identify concrete fake cues. While VBench and human-preference studies (Liu et al., 2025) provide holistic scores, they lack fine-grained localization. Our work instead gathers spatially and temporally grounded human annotations of perceived fakeness, offering a precise view of how humans judge generated videos.
Video Generation and Evaluations

### v2

@@ 1 Introduction
Recent advances in video generation technologies, including Veo3 (Anil et al., 2024), Sora (OpenAI, 2024), Pika (Pika, 2024), Meta Movie Gen (AI, 2024), Gen-3 (Runway Research, 2024), Kling (Kling, 2024), and others (Yang et al., 2024; Team, 2024; MiniMax, 2024; Wang et al., 2023a; Li et al., 2024b), have demonstrated remarkable capabilities in producing increasingly realistic videos. Alongside this progress, numerous studies about video generation have been conducted (Huang et al., 2023; Liu et al., 2023; Bansal et al., 2024; Huang et al., 2024; Liu et al., 2025), such as evaluating video prompt alignment towards a set of provided prompts as in VBench (Huang et al., 2023), or analyzing the physical commonsense in deepfake videos as explored by VideoPhy (Bansal et al., 2024), etc. However, these evaluations primarily compare AI-generated videos against a set of predetermined criteria, neglecting one of the most fundamental aspect:
@@ 4 Related Work
AI-generated Video Detection
Prior research has extensively explored the detection of machine-generated text (Dugan et al., 2024; Ippolito et al., 2020) and AI-generated images (Guo et al., 2023; Lorenz et al., 2023; Wu et al., 2023; Wang et al., 2023b). With the emergence of video generation models, several techniques have also been proposed for fake video detection (Gu et al., 2022; Xu et al., 2024; Gu et al., 2021). However, most of these efforts are narrowly focused on human face deepfakes, whereas current video generation models are capable of producing much more diverse and complex content. Recent work such as Beyond the Imitation Game (Vahdati et al., 2024) attempts to detect synthetic videos by learning generalized traces of video generation artifacts. Similarly, DeMamba (Chen et al., 2024) introduces the large-scale GenVideo dataset, comprising both fake and real videos, and improves detection performance by leveraging spatiotemporal inconsistencies. Despite these advancements, both methods treat fake video detection as a binary classification problem and do not require models to explain or localize the underlying reasons for fakeness. In contrast, our work goes a step further by collecting fine-grained human rewards that identify not only whether a video is fake, but also where and why the fakeness occurs.